OpenStudy (anonymous):

Let f(x) = lx-1l. Which of the following conditional statements hold true? *Check all that apply* A. f(x) = x - 1 if x > 1 B. f(x) < 0 if x = 0 C. f(x) = x - 1 if x > 0 D. f(x) = -(x - 1) if x < 1

OpenStudy (anonymous):

do you know the definition of \(|x-1|\) ?

OpenStudy (anonymous):

i mean the real math definition, not the "make it positive" definition

OpenStudy (anonymous):

No. But an explanation would be much appreciated :D

OpenStudy (anonymous):

ok lets start with the real definition of \(f(x)=|x|\) which is \[f(x) = |x| = \left\{\begin{array}{rcc} -x & \text{if} & x <0 \\ x& \text{if} & x \geq 0 \end{array} \right. \]

OpenStudy (anonymous):

a piecewise function reads \(|x|\) is \(-x\) if \(x\) is negative (which of course makes \(-x\) positive) or just \(x\) if \(x\) is positive (since it is already positive, don't change it)

OpenStudy (anonymous):

now we replace \(x\) by \(x-1\) to get your example \[f(x) = |x-4| = \left\{\begin{array}{rcc} -(x-1) & \text{if} & x-1 <0 \\ x-1& \text{if} & x-1 \geq 0 \end{array} \right. \]

OpenStudy (anonymous):

we really want to write this as \[f(x) = |x-4| = \left\{\begin{array}{rcc} -(x-1) & \text{if} & x <1 \\ x-1& \text{if} & x \geq 1 \end{array} \right.\]

OpenStudy (anonymous):

in other words the absolute value of \(x-1\) is just \(x-1\) so long as \(x\geq 1\) i.e. as long as \(x-1\) is positive, but if \(x-1\) is negative then \(|x-1|\) is \(-(x-1)\)

OpenStudy (anonymous):

now we can check all that apply \(A. f(x) = x - 1 \) if \( x > 1\) YES for sure that is the second line of the definition

OpenStudy (anonymous):

B is just plain silly, the absolute value is never negative, so NO

OpenStudy (anonymous):

how about C?

OpenStudy (anonymous):

I think C applies

OpenStudy (anonymous):

Because x is greater then 0

OpenStudy (anonymous):

?